OpenStudy (anonymous):

can someone please help me with Cramer's rule for two-variable systems? Here are the two equations: 3x-2y=7 2x+7y=38

OpenStudy (anonymous):

3 -2 2 7 get the determinant, this will be |A| then replace the 7 into 3 so it will be 7 -2 38 2 38 7. This will be |Ax|. x = |Ax|/|A| do this for y

OpenStudy (jdoe0001):

\(\large {\begin{bmatrix} 3&-2&|&\color{red}{7}\\ 2&7&|&\color{red}{38} \end{bmatrix}\\ \quad \\ D=\begin{bmatrix} 3&-2\\ 2&7 \end{bmatrix}\qquad D_x = \begin{bmatrix} \color{red}{7}&-2\\ \color{red}{38}&7 \end{bmatrix}\qquad D_y= \begin{bmatrix} 3&\color{red}{7}\\ 2&\color{red}{38} \end{bmatrix}\\ \quad \\ x = \cfrac{\textit{determinant of }\begin{bmatrix} \color{red}{7}&-2\\ \color{red}{38}&7 \end{bmatrix}}{\textit{determinant of }D}\qquad y = \cfrac{\textit{determinant of }\begin{bmatrix} 3&\color{red}{7}\\ 2&\color{red}{38} \end{bmatrix}}{\textit{determinant of }D}}\)